OpenStudy (anonymous):

FAN AND MEDAL AND TESTIMONIAL!! art A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

all x's are exponents...

OpenStudy (anonymous):

is part A because they both are y=something so you can equate them together?

OpenStudy (princeharryyy):

solve equation => y=8x and y=2x +2 which gives x= 1/3... which satisfies the equation 8x = 2x +2

OpenStudy (princeharryyy):

hence the solution

OpenStudy (anonymous):

@camerondoherty @dan815

OpenStudy (princeharryyy):

i think u can make the table urself

OpenStudy (anonymous):

dan please say you can help explain part A?

OpenStudy (princeharryyy):



OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

they arent linear they are exponents.

OpenStudy (princeharryyy):

ohk den

OpenStudy (princeharryyy):

no problem process still remains the same

OpenStudy (anonymous):

i need part A: not the answer

OpenStudy (princeharryyy):

ohk

OpenStudy (anonymous):

@dan815 help me please.

OpenStudy (dan815):

because at that point is where the x,y values of both eqations are the same

OpenStudy (dan815):

so that x value when input into both equations yieds the same y

OpenStudy (anonymous):

well idk part a

OpenStudy (princeharryyy):

when u equate the two equations => u get this 8^x = 2^x +2 which is same as the equation that u r given ... that is 8^x = 2^x +2 this. hence solution of the curve satisfies the equation @knightmare6

OpenStudy (anonymous):

DUDE I DONT GET WHAT YOU MEAN I JUST NEED PART A TO CONTINUE

OpenStudy (anonymous):

dan. can you please explain

OpenStudy (anonymous):

@dan815

OpenStudy (princeharryyy):

that is part A @knightmare6

OpenStudy (anonymous):

okay now to part B thanks

OpenStudy (anonymous):

that cant be because it says something about the x cordinate.

OpenStudy (princeharryyy):

graph 1 => y = 8^x graph 2 => y = 2^x +2 plug in different x's between -3 and 3 .... and find different y's corresponding to those x's... when the y's become equal ... it will be the required solution .... that x is the solution.

OpenStudy (anonymous):

i got part A.

OpenStudy (princeharryyy):

@knightmare6 @ part b

OpenStudy (anonymous):

i did part B

OpenStudy (anonymous):

okay i finished all of it thanks for trying to help me out i just got confused.

OpenStudy (princeharryyy):

:)